# Matryoshka: Non-Exclusive Memory Tiering via Transactional Page Migration

### v1

@@ 4 Evaluation
Workload type Success : Aborted
10GB WSS: read 148 : 1
10GB WSS: write 197 : 1
13.5GB WSS: read 151 : 1
13.5GB WSS: write 267 : 1
27GB WSS: read 167 : 1
27GB WSS: write 153 : 1
Table 3: The success rate of transactional migration.
Experimental setup: We compared Matryoshka with two state-of-the-art tired memory systems: TPP [41] and Memtis [34]. We conducted experiments upon three platforms with different configurations in CPU, local DRAM, CXL memory, and persistent memory, as listed in Table 1. Notably, we used a real-world, Agilex-7 FPGA-based CXL memory device [6] and Intel Optane 100 Series Persistent Memory (PM). The CXL memory device has only 16 GB of memory. Hence, we configured local DRAM with a size of 16 GB  [^3] Owing to the limited availability of CXL memory devices, we could not test Matryoshka with a larger CXL memory tier. However, Matryoshka’s non-exclusive design would benefit more from a large-capacity slow tier.. The size of the Intel Optane PM was 256 GB. In this tiered memory, local DRAM serves as the fast memory tier, while CXL or PM serves as the slow memory tier. In addition, each of the three platforms was equipped with a single socket, and we configured the number of CPU cores to 32. We evaluate both TPP and Matryoshka on Linux kernel v5.15, and run Memtis on kernel v5.19. We first used micro-benchmarks to investigate the performance of Matryoshka’s transactional page migration and page shadowing mechanisms. Further, we used real-world applications to assess the overall performance of Matryoshka.
@@ 4.1 Micro-benchmarks
We developed a micro-benchmark to precisely assess Matryoshka in a controlled manner. This micro-benchmark involves 1) allocating data to specific segments of the tiered memory, 2) running with various working set sizes (WSS) and resident set sizes (RSS); and 3) generating memory accesses to the WSS data that mimic real-world memory access patterns with a Zipfian distribution. We created three distinct scenarios, which represent the small, medium, and large WSS, for the micro-benchmarking. We conducted all the micro-benchmarking on Platform B – a tiered memory system with 16 GB local DRAM and 16 GB remote CXL memory.
@@
A small WSS: We began with a scenario by running the micro-benchmark with a small WSS of 10 GB and a total RSS of 20 GB. Initially, we filled the first 10 GB of local DRAM with the first half of the RSS data. Subsequently, we allocated 10 GB of WSS data as the second half of the RSS data: 6 GB on the local DRAM and 4 GB on the CXL memory. The micro-benchmark, during its operation, continuously executed memory reads or writes (following a Zipfian distribution) to this 10 GB WSS data, spread across both the local DRAM and CXL memory. The frequently accessed data, or “hot” data, was uniformly distributed along the 10 GB WSS. For both TPP and Matryoshka, accessing data on the CXL memory triggered its migration to the local DRAM, with TPP doing so synchronously and Matryoshka asynchronously. In contrast, Memtis used a background thread to migrate hot data from the CXL memory to the local DRAM. Due to page migration, the 4 GB data on the CXL memory gradually moved to the local DRAM – i.e., a part of the cold RSS data was demoted to the CXL memory. As the total WSS was small (i.e., 10 GB), it was stored in the local DRAM after the micro-benchmark ran for a while, reaching a stable state.
Figure 7 illustrates that in the initial phase, where intensive page migration was conducted (i.e., migration in progress), both Matryoshka and Memtis demonstrated equivalent performance regarding memory bandwidth for reads and writes. During this period, they outperformed TSS significantly, with improvements ranging from 100% to 300%. The performance benefit came from the asynchronous, transactional page migration for Matryoshka (Section 3.1) and asynchronous, background page migration for Memtis, which merely blocked the execution of the micro-benchmark.
@@
Figure 7 further shows the results in the stable phase (i.e., migration stable), where most of the WSS data was migrated from the CXL memory to the local DRAM. Both Matryoshka and TPP achieved equivalent read/write performance because memory accesses were all served by the local DRAM and there were few or no page migrations, as listed in Table 2. Surprisingly, the performance of Memtis was the worst (less than 40% of the other two). In fact, the performance with Memtis did not change much from the initial phase to the stable phase. Possible reasons include 1) Memtis’s PEBS-based sampling was too coarse-grained and failed to capture most hot pages on the CXL memory; 2) Memtis’s background migration thread could not timely migrate pages from the local DRAM to the CXL memory. We only observed a small number of page migrations occurring during both the initial and stable phases (not listed in Table 2). Hence, in the stable phase, many memory accesses were still served by the slow CXL memory, leading to low performance.
A medium WSS: We continued to increase the WSS and the RSS to a larger size, setting both at 13.5 GB and 27 GB separately. Once again, we placed the first half of the RSS data (13.5 GB) at the start of the local DRAM, followed by 2.5 GB of the WSS on the local DRAM, with the remaining 11 GB residing on the CXL memory.
Similar to the small WSS case, Figure 8 demonstrates that during the initial phase, where heavy page migrations were conducted (migration in progress), both Matryoshka and Memtis achieved comparable memory bandwidths for read and write operations. In this period, they significantly outperformed TPP, showing improvements in the range of 2x to 3x. Compared with the small WSS case, in the medium WSS case, more page migrations were conducted in the initial phase – e.g., more than 2x as listed in Table 2. This is because a larger portion of the WSS data was initially allocated in the CXL memory than the small WSS case (11 GB vs. 4 GB).
Figure 8 presents the findings from the stable phase (migration stable). In this phase, Matryoshka and Memtis exhibited nearly the same performance in terms of write operations. However, for read operations, Matryoshka notably excelled, outperforming the other two approaches substantially – for instance, it achieved 4x the performance of TSS and 2x the performance of Memtis. In the medium WSS case, we observed that, during the stable phase, a significant number of page migrations were still triggered, as listed in Table 2. The reason is that although the WSS of 13.5 GB is smaller than the local DRAM capacity of 16 GB, the combined size of the WSS data and the system’s data exceeds the threshold for space reclaiming, triggering the demotion thread to begin migrating pages to the CXL memory. The demoted pages belonging to the WSS were promptly promoted back to the local DRAM shortly, leading to memory thrashing.
Matryoshka can significantly mitigate the performance degradation for read operations (resulting from memory thrashing) due to its non-exclusive memory tiering. Recall that Matryoshka’s page shadowing mechanism enables a subset of pages resident in the local DRAM to have a shadow copy in the CXL (Section 3.2). This mechanism benefits the read operations when thrashing occurs – when a page is selected for demotion and meanwhile Matryoshka retains its copy in the CXL memory, the page is discarded rather than migrated. Unfortunately for writes, as the local version has been modified during demotion, the new version needs to be migrated to the CXL memory. But Matryoshka still managed to deliver the best performance among the three approaches.
@@
A large WSS: In our final test using micro-benchmarks, we scaled up the WSS (and the RSS) to a size of 27 GB. In this scenario, all the data accessed in the micro-benchmark constituted the WSS data. We arranged this data sequentially, starting from the local DRAM and extending to the CXL memory, allocating 16 GB of the WSS on the local DRAM and the remaining 11 GB on the CXL memory. This created an extreme scenario that caused severe memory thrashing, as the total hot data greatly (27 GB) exceeded the capacity of the local DRAM (16 GB).
@@
Figure 9 presents the performance in the initial phase and the the stable phase. Interestingly, both Matryoshka and TPP, tending to promptly migrate detected hot data from the CXL memory, performed worse than Memtis, which only conducted background promotion/demotion. This is because, the severe memory thrashing caused by Matryoshka and TPP, frequently migrated pages between the local and remote memory, greatly degrading the micro-benchamrk’s performance. In contrast, much fewer page migrations occurred with Memtis. However, Matryoshka consistently outperformed TPP by a range of 2x - 5x, again due to its asynchronous, transactional page migration. In addition, Matryoshka’s page shadowing helped achieve higher performance. Regardless, frequent page migrations become counterproductive when the WSS significantly exceeds the capacity of fast memory.
@@
Migration success rate: As discussed in Section 3.1, there exists a possibility that Matryoshka’s transactional page migration might not succeed due to simultaneous modification (dirtying) of the page by a process. Table 3 demonstrates that, across various scenarios tested with micro-benchmarks, Matryoshka maintains a notably high success rate in page migrations.
@@
Table 4: Robustness evaluation on Platform B with 30.7GB tiered memory.
Robustness: Page shadowing can potentially increase memory usage and in the worst case can cause OOM errors. In this test, we evaluated the effectiveness of Matryoshka’s shadow page reclamation. We measured the total memory usage and the size of shadow memory using a micro-benchmark that sequentially scans a predefined RSS area. Table 4 shows the change of shadow pages as we varied the RSS. The results suggest that Matryoshka was able to effectively reclaim shadow pages as the benchmark approached the capacity of the total memory. We further pushed the RSS (not shown in the table) beyond the total memory size. The system started swapping but encountered no OOM errors. We repeated the tests for a large number of times and no OOM errors were observed.
Summary: Matryoshka enables applications to sustain high performance via asynchronous, transactional page migrations when the WSS is smaller than the performance tier. Additionally, when the WSS is comparable to or exceeds the capacity of the performance tier, Matryoshka’s page shadowing feature notably enhances the efficiency of read operations.
Figure 10: Performance comparisons using Redis and YCSB between TPP and Matryoshka. Despite our best efforts, we could not execute Memtis on Redis successfully.
Figure 11: Performance comparisons with PageRank between non-migration, TPP, Memtis, and Matryoshka.
Figure 12: Performance comparisons with Liblinear between non-migration, TPP, Memtis, and Matryoshka.

### v2

@@ 4 Evaluation
Workload type Success : Aborted 10GB WSS: read 148 : 110GB WSS: write 197 : 1 13.5GB WSS: read 151 : 1 13.5GB WSS: write 267 : 1 27GB WSS: read 167 : 1 27GB WSS: write 153 : 1 The success rate of transactional migration.
Workload type Success : Aborted Liblinear (large RSS) on platform C 1:1.9 Liblinear (large RSS) on platform D 2.6:1 Redis (large RSS) on platform C 153:1 Redis (large RSS) on platform D 278.2:1 Success rate of Nomad’s migration (applications).
Testbeds. We conducted experiments upon on four platforms with different configurations in CPU, local DRAM, CXL memory, and persistent memory, as detailed in Table 1.
  • Platform A was built with commercial off-the-shelf (COTS) Intel Sapphire Rapids processors and a 16 GB Agelix-7 Agilex-7 FPGA-based CXL memory device [agilex].
  • Platform B featured an engineering sample of the Intel Sapphire Rapids processors with the same FPGA-based CXL memory device. The prototype processors have engineering tweaks that have the potential to enhance the performance of CXL memory, which were not available on Platform platform A.
  • Platform C included an Intel Alder Cascade Lake processor and six 128 256 GB 100 series Intel Optane Persistent Memory. This platform enabled the full tracking capability of Memtis capability of PEBS-based memory tracking and allowed for a comprehensive comparison between page fault- and sampling-based page migrationapproaches.
  • Platform D had an AMD Genoa 9634 processor and four 256 GB Micron’s (pre-market) CXL memory modules. This platform allowed us to evaluate Nomad with more realistic CXL memory configurations.
Since the FPGA-based CXL memory device had only 16 GB of memory, we configured local DRAM to 16 GB for all platforms  [^3] Although Platform platform C and D have larger PM or CXL memory sizes, we configured them with the same 16 GB DRAM to be slow memory consistent with Platform platform A and B for ease of a fair comparison in micro-benchmarks.This limit was lifted when testing real applications.. Note that Platform C was installed platform C was equipped with DDR4 (i.e., compatible with older processors), DRAM as fast memory while the other platforms were installed with used DDR5 . In this tiered memory , local DRAM served as the fast memory tier, while CXL or PMserved as the slow memorytier. In addition, each of the four platforms was equipped with a single socket, and we configured the number of CPU cores to 32 for Intel processors and 84 for AMDDRAM. We evaluated both CXL memory and persistent memory (PM) as slow memory. Table 1 also lists the main performance characteristics for the tiered memory of different platforms , including memory access latency and throughput (under lists the performance characteristics of the four platforms for single-threaded and peak performance) . (multi-threaded) performance. While CXL memory and PM have distinct characteristics, including persistence, concurrent performance, and read/write asymmetry, they achieve comparable performance within 2-3x of DRAM and provide a similar programming interface as a CPUless memory node. To ensure a fair comparison, we only enabled one socket on each of the four platforms. Intel platforms were configured with 32 cores while the AMD platform had 84 cores.
Baselines for comparison. We compared Nomad with two state-of-the-art tired memory systems: TPP [Maruf_23asplos_tpp] and Memtis [memtis]. We evaluated both TPP and Nomad on Linux kernel v5.13-rc6 and ran Memtis on kernel v5.15.19, the kernel version upon which Memtis was built and released. We tested two versions of Memtis – Memtis-Default and Memtis-QuickCool – with different data cooling speeds (i.e., the period to halve number of samples collected before halving a page’s access countercount). Specifically, Memtis-Default chose used the default cooling period (of 2,000,000)000k samples, while Memtis-QuickCool was configured with a much smaller cooling period (of 2,000)used a period of 2k samples. A shorter cooling period instructs Memtis to retain fewer pages in the hot list and demote pages more quickly when the fast tier is under memory pressureencourages more frequent page migration between the memory tiers.
Note that
[…]
(a) Small working set size
[…]
(b) Medium working set size
[…]
(c) Large working set size
Figure 7: Performance comparison between TPP, Memtis-Default, Memtis-QuickCool, and Nomad on platform A.
Memtis relies on Intel’s Processor Event-Based Sampling (PEBS) to track memory access patterns. Using PEBS, Memtis It samples various hardware events, such as including LLC misses, TLB misses, and retired store instructions. By sampling these events, Memtis can detect which pages (from the fast or slow tier) are accessed by walking through the process’s page table. We observed two facts with Memtis : 1) It does not work on Platform D, which is equipped with the AMD processor – although AMD has a PEBS-equivalent feature called Instruction-Based , to infer accessed page addresses and build frequency-based histograms to aid in making migration decisions. Memtis currently only supports Intel-based systems, though it can be ported to AMD processors with Instruction-based Sampling (IBS), Memtis does not support it. Hence, we did not evaluate Memtis on Platform D. 2) . Thus, Memtis was not evaluated on platform D. Memtis works slightly differently between on CXL-memory platforms (Platform systems (platforms A and B) and PM platforms (Platform the PM system (platform C). It is because some hardware events related to CXL memory, e.g., LLC misses to CXL memory , are not supported by PEBS and are regarded as uncore events on Intel platforms and thus cannot be captured by Memtis. As a resultPEBS. Therefore, Memtis relies solely on TLB misses and retired store instructions to track CXL memory access patterns. However, we also noticed that too many LLC misses to local DRAM may overshadow TLB misses stemming from accessing CXL memory, reducing the accuracy of memory access tracking. infer page temperature on platforms A and B.
@@ 4.1 Micro-benchmarks
We first used micro-benchmarks to investigate To evaluate the performance of Nomad’s transactional page migration and page shadowing mechanisms. To do thisshadowing mechanisms, we developed a micro-benchmark to precisely assess Nomad in a controlled manner. This micro-benchmark involves 1) allocating data to specific segments of the tiered memory, ; 2) running tests with various working set sizes (WSS) and resident set sizes (RSS); and 3) generating memory accesses to the WSS data that mimic real-world memory access patterns with a Zipfian distribution. We created three distinct scenarios representing small, medium, and large WSSfor the micro-benchmarking. We conducted all the micro-benchmarking on all platforms. Note that Platform B produced similar results as A (hence not listed). , as illustrated in Figure 6, to evaluate tiered memory management under different memory pressures. As platform B behaved similarly to platform A in micro-benchmarks, it is excluded from the discussion.
@@
Small WSS. We began with a scenario by running the micro-benchmark with a small WSS of 10 GB and a total RSS of 20 GB. Initially, we filled the first 10 GB of local DRAM with the first half of the RSS data. Subsequently, we allocated 10 GB of WSS data as the second half of the RSS data: 6 GB on the local DRAM and 4 GB on the CXL memory (Platform platforms A and D) and the PM (Platform platform C). The micro-benchmark continuously performed memory reads or writes (following a Zipfian distribution) to this 10 GB WSS data, spread across both the local DRAM and CXL memory or PM. The frequently accesseddata, or “hot” data, was uniformly distributed along the 10 GB WSS. For both In TPP and Nomad, accessing data on the CXL memory or PM triggered its page migration to the local DRAM, with TPP performing this migration synchronously and Nomad asynchronously. In contrast, Memtis used a background thread to migrate hot data from the CXL memory or PM to the local DRAM. Due to page migration, the 4 GB data on the WSS data, initially allocated to CXL memory or PM, was gradually moved to the local DRAM, while part of the cold RSS data was demoted to the CXL memory or PM. Since the WSS was small (i.e., 10 GB), it would could be completely stored in the fast tier (i.e., local DRAM) after the micro-benchmark ran for a while, reaching reached a stable state.
[…]
@@
In the the stablephasestable phase (i.e., migration stable), where migration stable), when most of the WSS data had been migrated from the CXL memory or PM to the local DRAM, both Nomad and TPP achieved similar read/write bandwidth. This was because memory accesses were primarily served by the local DRAM , with few page migrationsoccurring, as shown in Table 2. Memtis performed the worst, achieving as low as 40% of the performance of the other two approaches. We further observed that, in most cases, We make two observations regarding Memtis’s performance increased only slightly from the initial to the stable phase. Furthermore, with a higher rate of cooling down hot pages, Memtis-QuickCool performed better than Memtis-Default. Possible reasons for Memtis’s low performance include: 1) Memtis’s PEBS-based sampling was too coarse-grained and failed to capture most hot pages on the CXL memory (i.e., as stated earlier, LLC misses to CXL memory cannot be captured by Memtis); and 2) Memtis’s background migration thread could not migrate pages from the local DRAM to the CXL memory promptly. We observed only a small number of page migrations conducted by Memtis during both the initial and stable phases, as listed weaknesses. First, its stable phase performance is not drastically different from the transient phase. The migration statistics in Table 2 . Consequently, show that Memtis performed significantly fewer page migrations. This explains its sub-optimal performance in the stable phase , the slow memory tier still handled many memoryaccesses, leading to low performance. as most memory accesses were still served from slow memory. Second, a shorter cooling period in Memtis, which incentivizes more frequent migrations, led to better performance. This also suggests that sampling-based memory access tracking may not accurately identify and timely migrate hot pages to fast memory.
We continued to increase the WSS and the RSS to larger sizes, setting them at
[…]
(a) Small working set size
[…]
(b) Medium working set size.
[…]
(c) Large working set size
Figure 8: Performance comparison between TPP, Memtis-Default, Memtis-QuickCool, and Nomad on platform C.
[…]
(a) Small working set size
[…]
(b) Medium working set size
[…]
(c) Large working set size
Figure 9: Performance comparison between TPP and Nomad on platform D with AMD Genoa processor. Memtis does not support AMD’s instruction-based sampling (IBS) and thus was not tested.
Medium WSS. We increased the size of WSS and RSS to 13.5 GB and 27 GB, respectively. Once againSimilarly, we placed the first half of the RSS data (13.5 GB) at the start of the local DRAM, followed by 2.5 GB of the WSS on the local DRAM, with the remaining 11 GB residing on the CXL memory or PM. However, as the system (e.g., the OS kernel) required approximately 3-4 GB of memory, the entire WSS could not be fully held WSS could barely fit in the fast tierduring the stable state, resulting in constant page migrations . occasional and substantial migrations even during the stable phase. Accurately identifying hot pages and avoiding thrashing is crucial to achieving high performance for this medium-sized benchmark.
Unlike the small WSS case, Figures 7 (b), 8 (b), and 9 (b) show that during the initial transient phase, Nomad and TPP generally achieved lower memory bandwidth for performance for both read and write operations compared to Memtis. This is because, under the medium WSS, the system experienced higher memory pressure than in the small WSS case, causing Nomad and TPP to conduct more page migrations (2x - 6x) and incur higher overhead than Memtis, as shown in Table  2. Many of such migrations were futile during thrashing. Conversely, Memtis performed significantly fewer page migrations than the small WSS case. One possible reason is that : It did not capture the LLC misses of CXL memory and mostly relied on TLB misses for tracking memory access, but the events of TLB misses were overshadowed by extensive local LLC miss events in the medium WSS case. The less effective PEBS-based approach in Memtis, resulting in fewer page migrations, redemptively maintained the system’s and avoided the waste. However, there was no evidence that Memtis effectively detected thrashing and throttled migration. The coarse-grained sampling was unable to accurately determine page temperature in a volatile situation and inadvertently sustained high performance under high memory pressure.
As the system entered
In the stable phase, Nomad achieved higher performance for both reads and writes than the initial phase due to 1)most hot pages residing in the fast tier and thus 2) fewer page migrations being triggered. Compared to Memtis-Default, Nomad achieved higher read performance and similar write performance . The reason that Nomad was particularly superior in read performance – the read bandwidth was the highest among all test cases during the stable state across all three platforms – was due to significantly outperformed TPP in all cases, especially on platform D. These results show the benefit of Nomad’s ’s transactional page migration and non-exclusive memory tiering. Recall that memory tiering compared to TPP’s synchronous page migration and exclusive tiering. On platform D, which was equipped with an application-specific integrated circuit (ASIC)-based CXL memory implementation, the performance gap between fast and slow memory narrows. Thus, the software overhead associated with synchronous page migration was exacerbated and Nomad ’s page shadowing mechanism enables a subset of pages resident in the local DRAM to have a shadow copy in the CXL memory or PM (Section 3.2). This mechanism benefits the read operations when thrashing occurs (e.g., with a medium WSS ) – when a page is selected for demotion and meanwhile offered more pronounced performance gains. Similarly, Nomad retains its copy in the CXL memory or PM, achieved substantially higher performance in reads than Memtis and comparable performance in writes. Unlike in the page is discarded rather than migratedsmall WSS case, in which asynchronous and transactional page migration in Nomad contributed most to its performance benefit, the advantage of page shadowing played a critical role in alleviating thrashing in the medium WSS case. Under memory thrashing, most demoted pages, which were recently promoted from the slow tier, can be simply discarded without migration. However, as discussed earlier, the page shadowing mechanism could be detrimental to write operations due to additional pagefaultsfor write-intensive workloads, page shadowing requires one additional page fault for each write to restore a page’s original read-write permission. This explains why Nomad achieved similar write performance as Memtis-Default and sometimes worse than Memtis-QuickCool. Finally, Nomadagain significantly outperformed TPP in all cases. ’s inferior write performance in the stable phase compared to Memtis.
@@
Large WSS. We scaled up the WSS (and RSS ) and RSS both to 27 GB and arranged the data from the local DRAM to the CXL memory or PM, with fully populated local DRAM with the first 16 GB of the WSSon the local DRAM and the remaining on the . The remaining WSS spilled onto CXL memory or PM. This created an extreme case causing severe Unlike the medium WSS that incurred intermittent memory thrashing, as the total this workload caused continuous and severe thrashing as the size of hot data greatly exceeded the capacity of the local DRAM. fast memory.
(a) Small working set size (b) Medium working set size (c) Large working set size Performance comparison between TPP, Memtis-Default, Memtis-QuickCool, and Nomad on platform A.
(a) Small working set size (b) Medium working set size. (c) Large working set size Performance comparison between TPP, Memtis-Default, Memtis-QuickCool, and Nomad on platform C.
(a) Small working set size (b) Medium working set size (c) Large working set size Performance comparison between TPP and Nomad on platform D with AMD Genoa processor. Memtis does not support AMD’s instruction-based sampling (IBS) and was not tested.
@@
Figures 7 (c), 8 (c), and 9 (c) present the performance results in both the initial phasetransient phase and the the stablephasestable . As expectedphase. Compared to the tests with the medium-sized workload in which Nomad could outperform Memtis for read-only benchmarks, especially in the stable phase, both Nomad and TPP , which tend to promptly migrate detected hot data from the CXL memory or PM, performed worse than Memtis , which only conducted background promotion/demotion. This is because the in almost all scenarios. It suggests that page fault-based tiered memory management, which makes per-page migration decisions upon access to a page, inevitably incurs high overhead during severe memory thrashingcaused by the large WSS made Nomad and TPP frequently migrate pages between the local and remote memory, greatly degrading the micro-benchmark’s performance. In contrast, far fewer page migrations occurred with Memtis. Additionally. Nevertheless, Nomad consistently outperformed TPP due to its and significantly outperformed TPP thanks to asynchronous, transactional page migration . Again, Nomad’s page shadowinghelped achieve higher read performance, especially during the stable state on Platform A. Regardless, frequent page migrations become counterproductive when the WSS significantly exceeds the capacity of fast memory. page migration and page shadowing.
Figure 10: A PEBS-tailored scenario The average cache line access latency on Platform platform C. The memory was accessed in blocks of 1GB, within which the workload followed benchmark is a pointer-chasing point-chasing workload optimized for PEBS-based memory access patterntracking.
Our observations from the above test cases showed that Memtis performed page migrations much less frequently than Nomad and TPP. In the final micro-benchmarking test, we created a PEBS-tailored scenario designed to generate a PEBS event for each memory access, potentially allowing Limitations of PEBS-based approaches. Our evaluation revealed several issues with PEBS-based Memtis to capture memory access more accurately and result in more page migrations. We used a random pointer-chasing memory access pattern, where memory access tracking. While Memtis prevented excessive migrations during thrashing, it achieved sub-optimal performance and failed to migrate all hot data to fast memory even when the WSS could fit in the fast tier. Due to the lack of hardware support for hot page tracking, PEBS-based approaches employ indirect metrics, such as LLC and TLB misses to sample recently accessed addresses to infer page temperature. Sampling-based memory tracking has two fundamental limitations. First, there is a difficult tradeoff between sampling rate and tracking accuracy. Second and most importantly, cache misses may not effectively capture hot pages. For most frequently accessed pages that always hit the caches, Memtis fails to collect enough (cache miss) samples to build the histogram. If such pages are evicted from the caches, e.g., due to conflict or coherence misses, they will be falsely regarded as “cold” pages.
To demonstrate these limitations, we created a favorable scenario where Memtis can capture every page access. We used a pointer-chasing benchmark that repeatedly accesses multiple fixed-sized (1 ) each memory access resulted in a cache miss (a PEBS event on Platform C equipped with PM) and 2) access to memory blocks (at the granularity of GB) memory blocks. Within each 1 GB) exhibited temporal locality. That is, each page would be accessed multiple times (4KB / 64B = 64 times) within a memory block, and the same memory block might be accessed repeatedly in the future. This way, accurate page migration would significantly benefit the workload. GB block, the benchmark randomly accesses all cache lines belonging to a block while accesses across blocks follow a Zipfian distribution. The number of blocks determines the WSS. Since the block size exceeds the LLC size in our testbeds, every access generates an LLC miss that can be captured by Memtis. Effective memory access tracking should identify hot blocks and place them in fast memory.
Figure 10 shows that, under the medium and large WSS on Platform C, Memtis exhibited the slow-tier latency (i.e. , most pages in the average latency to access a cache line in this benchmark on platform C. Note that platform C with PM was the only testbed on which Memtis has full tracking capability and can capture all the PEBS events. According to Table 1, a latency closer to DRAM performance ($\sim$250 cycles) indicates more effective page placement. As shown in Figure 10, when the WSS exceeds fast tier capacity, Memtis achieved latency close to slow memory performance, suggesting that most hot pages still resided in the slow tier), while TPP and Nomad showed the fast-tier latency (i.e., most pages in the fast tier). It indicates that even with sufficient PEBS events, sampling-based PEBS Memtis still cannot accurately track memory access patterns and perform timely page migrations. . In comparison, page fault-based approaches, e.g., Nomad and TPP, can timely migrate hot pages and achieve low latency.
@@
As discussed in Section 3.1, there exists a possibility that Nomad’s transactional page migration might not succeed due to simultaneous modification (dirtying) of the page by a process. Table LABEL:tab:txSuccessRate demonstrates that, across various scenarios tested with micro-benchmarks, Nomad maintains a notably high success rate in page migrations.
@@
Table 3: Robustness evaluation Shadow memory size as RSS changes on Platform platform Bwith 30.7GB . The size of tiered memory (DRAM+CXL) is 30.7 GB.
Robustness:. Page shadowing can potentially increase memory usage and in the worst case can cause OOM errors if shadow pages are not timely reclaimed. In this test, we evaluated the effectiveness of Nomad’s shadow page reclamation. We measured the total memory usage and the size of shadow memory using a micro-benchmark that sequentially scans a predefined RSS area. Table 3 shows the change of shadow pages as we varied the RSS. The results suggest that Nomad was able to effectively reclaim shadow pages as the benchmark approached the capacity of the total memory . We further pushed the RSS (not shown in the table) beyond the total memory size. The system started swapping but encountered no OOM errors. We repeated the tests a large number of times and no OOM errors were observed.
When the WSS is smaller than the performance tier, Nomad enables workloads to sustain higher performance than Memtis through asynchronous, transactional page migrations. However, when the WSS is comparable to or exceeds the capacity of the performance tier, causing memory thrashing, page-fault-based migration in Nomad becomes detrimental to workload performance, falling short of Memtis for write operations. Notably, Nomad’s page shadowing feature retains the efficiency of read operations even in situations with severe memory thrashing, maintaining comparable or even better performance than Memtis in most cases. In all test cases, Nomad significantly outperformed the state-of-the-art page-fault-based migration approach, TPP. effectively reclaimed shadow pages to reduce shadow memory usage as RSS increased and approached memory capacity.